It was determined that the average weight of a newborn baby at the local hospital is 3.6 kg with a standard deviation of 0.7 kg. What is the probability that a newborn baby would weigh below 4.3 kg?


1
Expert's answer
2021-02-24T06:03:04-0500

"μ = 3.6 \\

σ = 0.7 \\

P(X<4.3) = P(\frac{x-μ}{σ} < \frac{4.3-3.6}{0.7}) \\

= P(Z < 1.0) \\

= 0.8413"

84.13 % is the probability that a newborn baby would weigh below 4.3 kg.